OpenStudy (bekkah323):

(working with absolute value). Solve. I3x+4I=Ix-7I

hartnn (hartnn):

if |a| = b then a= b or a=-b so, for |3x+4| = |x-7| you get, |3x+4| = x-7 or |3x+4| = -(x-7) which in turn , you get 3x+4 = x-7 or 3x+4 = -(x-7)

hartnn (hartnn):

can you solve those 2 equations to get x ?

OpenStudy (bekkah323):

the first one is 11/-2 = x and the second one is 3/4=x

OpenStudy (bekkah323):

@hartnn

hartnn (hartnn):

ok, x =- 11/2 and x= 3/4 are correct! good :)

OpenStudy (bekkah323):

is that it?

hartnn (hartnn):

yes, thats it :) quite simple, right ;)

OpenStudy (bekkah323):

yah, is that how all of those problems work?

hartnn (hartnn):

for = sign in between, all of the problems use if |a| = b then a= b or a=-b and work same way :)

OpenStudy (bekkah323):

ok thanks

hartnn (hartnn):

welcome ^_^